By how much does the sum of the roots of the equation ( x + 1) (x - 3) = 0 exceed the product of its roots?

Answer by mananth(16946)

(x+1)(x-3)
x=-1 and 3 are the roots of the equation
sum = -1+3= 2
product = 1*-3= -3
2-(-3)= 5
It exceeds by 5